# Multimodal Fusion and Coherence Modeling for Video Topic Segmentation

### v1

@@ 4.2 Pre-training with Unlabeled Data
Prior works have demonstrated that standard self-supervised denoising pre-training (even using only the downstream task data) Amos et al. (2023) or pre-training adapted to the downstream task Gururangan et al. (2020) often perform substantially better than randomly initializing the parameters. Therefore, to better initialize the parameters of the Multimodal Fusion Layers, we explore pre-training with unlabeled video data before fine-tuning on the labeled training set. Firstly, we introduce a general cross-modality alignment pre-training task to learn the multi-modal representation. We use contrastive learning loss to adjust the features learned by the model fusion layer, by maximizing the cosine similarity of the visual features and textual features of the same clip, while reducing the similarity of the modality features between different clips, as show in Eq. 11, where $\epsilon$ is used to prevent division by 0 and $\tau$ is a temperature hyper-parameter to scale the cosine similarity.
@@
Secondly, we introduce a novel pre-training task tailored for the downstream VTS task, focusing on utilizing unlabeled data for learning pseudo topic boundaries and also enhancing modality alignment. Initially, a Kernel Density Estimation (KDE) Davis et al. (2011) model is employed to approximate the topic duration distribution of the training set; then, videos are split into segments matching durations sampled from the KDE-derived distribution. Subsequently, for every resulting segment, we undertake one of three actions with an equal probability: inserting segments randomly selected from other videos either before or after the current segment, replacing the current segment with a segment randomly selected from other videos, or no change. The resulting segments within the modified video are then treated as distinct topics, thereby enabling the model to learn pseudo topic boundaries during pre-training. This task-adaptive pre-training task has the same $l_{vts}$ objective as shown in Eq. 5. The overall pre-training objective is shown in Eq. 13, where $\alpha$ and $\beta$ are hyper-parameters to adjust the loss weights.
@@ 4.3 Fine-tuning with Multimodal Coherence Modeling
For fine-tuning, we introduce two auxiliary tasks to enhance multimodal coherence modeling. The cross modal alignment task is the same task in Eq. 11 used in pre-training. This continuity ensures that the modalities retain their coherence through both pre-training and fine-tuning, fostering a consistent interplay between different modalities. In addition, we adapt the Contrastive Semantic Similarity Learning (CSSL) task proposed by Yu et al. (2023), which leverages the inherent characteristics of topic-related coherence, to the multimodal context. We adopt the same strategy for selecting positive and negative sample pairs Yu et al. (2023), but extend the features to the multimodal representations, as shown in Eq. 14, where $k_{1}$ and $k_{2}$ are hyper-parameters that determine the number of positive and negative pairs. For each clip’s multimodal representation $m_{i}$, $m_{i,j}^{+}$ denotes the multimodal representation of the $j$-th similar clip in the same topic as clip $i$, while $m_{i,j}^{-}$ denotes the multimodal representation of the $j$-th dissimilar clip in a different topic from clip $i$. We hypothesize that this extension could improve multimodal representation learning by identifying relative consistency relations within and across topics.
@@
The overall fine-tuning objective combines Eq. 5,  10,  11, and  14, as shown in Eq. 15, where $\sigma$, $\theta$, and $\gamma$ are hyper-parameters to adjust loss contribution. When the multimodl fusion layer does not contains MoE structure, $\beta$ in Eq. 13 and $\sigma$ are set to zero.
@@ 5.1 Experimental Setup
Datasets We partition the labeled data within AVLecture into 70% for training, 10% for validation, and 20% for test sets. We also split the CLVTS dataset into train, validation, and test sets with identical ratios. The unlabeled data of AVLectures and CLVTS are used for pre-training on each dataset, respectively.
Baselines We carefully select the following representative baselines:
- UnsupAVLS Gupta et al. (2023) is an unsupervised approach that clusters video clips into discrete topics based on visual and text embeddings learned from matching the narration with temporally aligned visual content. The actual number of topics is used as the desired number of clusters.
@@
- Text-only LongFormer is evaluated on long document topic segmentation by Yu et al. (2023). We fine-tune LongFormer and LongFormercssl in Yu et al. (2023) to evaluate the performance of a text-only model on the VTS task.
- Llama-3-8B [^6] https://huggingface.co/meta-llama/Meta-Llama-3-8B Our pre-training (Section 4.2) is conducted on the relatively limited unlabeled videos of AVLectures and CLVTS datasets. To investigate the effect of fine-tuning a powerful pre-trained text large language model (LLM) on VTS, we fine-tune Llama-3-8B with two prompting methods (see Appendix F for details).
- SWSTseq is our adapted version of the multimodal video scene and topic segmentation model Wang et al. (2023) to the sequence labeling VTS task, for comparing performance between their early fusion and our middle fusion strategy on VTS. SWSTseq uses a pre-trained LongFormer Beltagy et al. (2020) as backbone.
Evaluation Metrics We adopt four commonly used metrics for VTS, including positive F1 Zhang et al. (2021)(denoted as F1 to simplify notations), BS@k Gupta et al. (2023), mIoU Mun et al. (2022), and F1@k. Definitions of the four metrics are in Appendix D. We follow  Gupta et al. (2023) and set k to 30 seconds. Additionally, we use the average score (Avg) of these four metrics to measure the overall performance of a model.
Implementation Details We use the same maximum sequence length 2048 as  Yu et al. (2023) for a fair comparison with the text-only models. All results are the mean values over three runs with different random seeds. More implementation details are in Appendix C.
@@ 5.2 Results and Analysis
Table 2 reports the performance of baselines (the first group) and our MMVTS model variants (the second and third group) on the AVLecture and CLVTS test sets.
@@
Unimodal performance. The text-only Longformer (Row3) substantially outperforms the visual-only BaSSL (Row2) on BS@30 by a large gain (+25.31), and also surpasses the unsupervised approach UnsupAVLS by a notable gain (+13.25). Such results are expected since the textual modality inherently conveys more precise and richer information than the visual modality for VTS.
@@
Mutimodal performance. As can be seen from Table 2, Avg of $SWST_{seq}$ closely aligns with that of the text-only LongFormer, suggesting that more data may be necessary to mitigate the discrepancy between fine-tuning with early fusion and pre-training the language model, in order to fully exploit the potential of the early fusion strategy. Our MMVTS Baseline1,2,3 simply concatenate unimodal features to predict topic boundaries. Without pre-training on unlabeled data (PT, Eq. 13) and the two auxiliary fine-tuning tasks to enhance multimodal coherence modeling (FT-Coh, Eq. 15), on F1, MMVTS Baseline1 outperforms the text-only Longformer by 2.28 and 2.90 on AVLecture and CLVTS respectively, while on Avg score, MMVTS Baseline1 outperforms Longformer by 1.1 on AVLecture yet slightly underperforms Longformer by 0.65 on CLVTS. These results suggest that simply concatenating unimodal features to predict topic boundaries does not bring consistent gains over unimodality. The third group in Table 2 compares the four Multimodal Fusion Layer architectures with pre-training (PT) and fine-tuning (FT-Coh). Overall, compared to all the competitive unsupervised and supervised baselines, after pre-training and fine-tuning, our MMVTS model using Co-Attention with MoE as Multimodal Fusion Layers achieves the best Avg, BS@30, and F1@30 results and sets new SOTA on AVLecture, and achieves the best F1 and a Avg score comparable to the best result on CLVTS. The gains on Avg from our Co-Attention MoE with PT and FT-Coh over LongFormercssl and MMVTS Baseline1 are statistically significant (p-values are less than 0.05). It is also notable that the performance of models on CLVTS is generally much lower than that on AVLecture, with the best Avg score between AVLecture and CLVTS differing by 17.50 (68.84 versus 51.34), indicating a greater challenge posed by our CLVTS dataset than the AVLecture dataset.
@@
Table 2 also shows that on CLVTS, fine-tuning the powerful pre-trained Llama-3-8B with our Discrete prompt achieves the best Avg score 51.34, although very close to the performance of Merge-Attn with MoE and Co-Attn with MoE after PT and FT-Coh. However, on AVLecture, the performance of fine-tuning Llama-3-8B is much lower than Multimodal Fusion Layer with PT and FT-Coh; particularly, F1 on both AVLecture and CLVTS are much lower because the model’s predicted boundaries often have a clip offset, but its F1@30 and mIoU results are still best on CLVTS.
@@
Table 3: Ablation studies of the pre-training tasks on the AVLecture dataset. The two auxiliary coherence modeling tasks are added in fine-tuning (Eq. 15). The subscript of Avg score is its standard deviation.
We conduct ablation studies and analysis to validate effectiveness of the proposed multimodal fusion layers, pre-training and fine-tuning tasks.
@@
(2) Effect of Pre-training tasks. Table 2 shows that for simple concatenation of unimodal features, pre-training before fine-tuning (as Baseline3) outperforms Baseline2 (w/o PT). We conduct ablation studies of the proposed pre-training and fine-tuning tasks on the AVLecture dataset. We apply the same fine-tuning with multimodal coherence modeling (FT-Coh, Eq. 15) and compare (a) random initialization of parameters of Multimodal Fusion Layers (w/o pre-training) (b) pre-training the model on unlabeled data. Table 3 shows that w/o pre-training, Co-Attn slightly improves Avg over Merge-Attn by 0.4 pts, and MoE further improves Avg by 0.87 pts. Pre-training improves the performance on all four Multimodal Fusion Layer architectures, with the average Avg score increased by 1.36 pts (66.29 $\rightarrow$ 67.65). Pre-training also improves stability of a model as the standard deviations of all w/ PT experiments are less than 1. Table 6 in Appendix further compares the fine-tuning performance after applying different pre-training tasks. Compared to using both pre-training tasks, removing loss $l_{cma}$ or $l_{vts}$ leads to Avg reduction of 0.4 and 1.64 respectively, indicating that the pretraining task more closely aligned with the downstream task, i.e., $l_{vts}$, yields greater benefits.
@@
Table 4: Ablation studies of the two auxiliary coherence modeling fine-tuning tasks on the AVLecture dataset. Models are initialized from pre-training (Eq. 13).
(3) Effect of Fine-tuning tasks. Table 4 compares the efficacy of different fine-tuning tasks after pre-training. Compared to the standard $l_{vts}$, adding the two auxiliary losses $l_{cma}$ and $l_{mcssl}$ notably improves Avg by 1.12 pts, while decreases $F_{1}$ by 0.73. These results suggest that while multimodal coherence modeling may slightly compromise the precision of exact matches, it enhances the overall contextual comprehension of a model for VTS. Adding $l_{cma}$ or $l_{mcssl}$ individually improves Avg by 0.53 and 0.57 respectively, with improvements mainly on BS@30 and mIoU, suggesting that feature alignment at different granularities may improve fuzzy matching of a model.
@@ 6 Conclusion
In conclusion, we propose a supervised MultiModal Video Topic Segmentation (MMVTS) model to advances VTS by meticulously investigating multimodal fusion and coherence modeling, coupled with an innovative pre-training task tailored for VTS. Furthermore, we introduce a large-scale Chinese Lecture Video Topic Segmentation (CLVTS) dataset to promote research in VTS.
@@
It is essential to acknowledge that our methodology solely capitalizes on the multimodal information derived from visual and textual data, with the parameters of the visual encoder fixed due to consideration about computational complexity. This specific limitation potentially results in a less-than-optimal exploitation of the vast visual information contained within the dataset. Concurrently, features from additional modalities, such as audio, are not incorporated. Additionally, we will conduct more investigations and study approaches to make Co-Attn with MoE more stable in future work. Finally, the integration of general multimodal pre-trained models and large language models could be investigated, offering a more holistic and effective exploration of multimodal information for VTS.
Appendix A Related Work to VTS
Wu et al. (2023) utilized visual, audio and textual information, and focus on hierarchical modeling techniques to simultaneously learn the boundaries of scenes, stories and topics, which suggests that the segmentation can be benefited by introducing the hierarchical structures of videos. Nevertheless, their approach did not delve into optimizing the integration of multimodal information, which is one of the motivations of our work. The integration of audio features represents a direction for our future work. Xing et al. (2024) encode the image and text sequences with respective encoders and use an asymmetric cross-modal cross-attention mechanism to produce text-aware visual representations, followed by a BiLSTM, to fuse multimodal features. Our approaches are drastically different from Xing et al. (2024) as we investigate symmetric Cross-Modal Cross-Attention and also Mixture-of-Experts mechanism to fuse multimodal information. Xing et al. (2024) also investigate cross-modal and sentence level intra-modal contrastive learning to empower unsupervised domain adaptation, whereas we explore extending the topic level Contrastive Semantic Similarity Learning (CSSL) task proposed by Yu et al. (2023) to the multimodal setting.
Appendix B Details of Annotations for the CLVTS Dataset

### v2

@@ 3.2 Pre-training with Unlabeled Data
Prior works have demonstrated that standard self-supervised denoising pre-training (even only using the downstream task data) Amos et al. (2023) or pre-training adapted to the downstream task Gururangan et al. (2020) often perform substantially better than randomly initializing the parameters. Therefore, to better initialize the parameters of the Multimodal Fusion Layers, we explore pre-training with unlabeled video data before supervised fine-tuning. Firstly, we introduce a general cross-modality alignment pre-training task to learn the multimodal representation. We use contrastive learning loss to adjust the features learned by the Multimodal Fusion Layers, by maximizing the cosine similarity of the visual features and textual features of the same clip, while reducing the similarity of the modality features between different clips, as show in Eq. 11, where $\epsilon$ is used to prevent division by 0 and $\tau$ is a temperature hyper-parameter to scale the cosine similarity.
@@
Secondly, we introduce a novel pre-training task tailored for the VTS task, focusing on utilizing unlabeled data for learning pseudo topic boundaries and also enhancing modality alignment. We apply a Kernel Density Estimation (KDE) Davis et al. (2011) model to estimate the topic duration distribution within the labeled training set. Videos are segmented based on KDE-sampled durations. For each segment, with equal probability, we: insert a random segment from other videos, replace it with another, or retain it. These modified segments serve as distinct topics, allowing the model to learn pseudo topic boundaries during pre-training. This task-adaptive pre-training task has the same $l_{vts}$ objective as shown in Eq. 5. The overall pre-training objective is shown in Eq. 13, where $\alpha$ and $\beta$ are hyper-parameters to adjust the loss weights.
@@ 3.3 Fine-tuning with Multimodal Coherence Modeling
For fine-tuning, we introduce two auxiliary tasks to enhance multimodal coherence modeling. The cross modal alignment task is the same as the task in Eq. 11 used in pre-training. This continuity ensures that the modalities retain their coherence through both pre-training and fine-tuning stages, fostering a consistent interplay between different modalities. In addition, we adapt the Contrastive Semantic Similarity Learning (CSSL) task proposed by Yu et al. (2023), which leverages the inherent characteristics of topic-related coherence, to the multimodal context. We adopt the same strategy for selecting positive and negative sample pairs Yu et al. (2023), but extend the features to the multimodal representations, as shown in Eq. 14, where $k_{1}$ and $k_{2}$ are hyper-parameters that determine the number of positive and negative pairs. For each clip’s multimodal representation $m_{i}$, $m_{i,j}^{+}$ denotes the multimodal representation of the $j$-th similar clip in the same topic as clip $i$, while $m_{i,j}^{-}$ denotes the multimodal representation of the $j$-th dissimilar clip in a different topic from clip $i$. We hypothesize that this extension could improve multimodal representation learning by identifying relative consistency relations within topics and across topics.
@@
The overall fine-tuning objective combines Eq. 5,  10,  11, and  14, as shown in Eq. 15, where $\sigma$, $\theta$, and $\gamma$ are hyper-parameters to adjust loss contribution. When the Multimodl Fusion Layers do not contain MoE structure, $\beta$ in Eq. 13 and $\sigma$ are set to zero.
@@ 4.1 Experimental Setup
Datasets. Table 1 summarizes the statistics of various VTS datasets. It clearly shows that prior to our work, AVLecture Gupta et al. (2023) is the only publicly available large-scale labeled video dataset facilitating supervised VTS methods. To promote the research in VTS, we introduce a large-scale labeled Chinese Lecture Video Topic Segmentation dataset (CLVTS). Both AVLecture and CLVTS are sourced from educational videos, where VTS significantly enhances learning experiences. In terms of differences, in addition to the linguistic distinctness from the English lecture dataset AVLecture, CLVTS is characterized by its natural and uninterrupted long videos, a stark contrast to AVLecture, since nearly two-thirds of AVLecture are reassembled pre-segmented short videos. As shown in Table 1, CLVTS features a higher average number of topics per video than AVLecture. Details of the data collection and annotation procedure and analysis of the CLVTS dataset are in Appendix A. Importantly, we put careful ethical considerations for the datasets used in this research in Appendix A.3.
Baselines and Implementation Details. The implementation details are in Appendix B. We carefully select the following representative baselines.
- UnsupAVLS Gupta et al. (2023) is an unsupervised approach that clusters video clips into a pre-defined number of topics, based on visual and text embeddings learned from matching the narration with the temporally aligned visual content.
@@
- Text-only LongFormer is evaluated on long document topic segmentation by Yu et al. (2023). We fine-tune LongFormer and LongFormercssl in Yu et al. (2023) to evaluate the performance of a text-only model w/o and w/ Contrastive Semantic Similarity Learning (CSSL) on the VTS task.
- Llama-3-8B Our pre-training (Section 3.2) is conducted on the relatively limited unlabeled videos of AVLectures and CLVTS datasets. To investigate the effect of fine-tuning a powerful pre-trained text large language model (LLM) on VTS, we fine-tune Llama-3-8B [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B with 8B parameters, using two different prompts (see Appendix E for details).
- SWSTseq is our adapted version of the multimodal video scene and topic segmentation model Wang et al. (2023) with a pre-trained LongFormer Beltagy et al. (2020) as the backbone to VTS, for comparing performance between their early fusion and our middle fusion strategy on VTS.
In this work, we choose not to utilize pre-trained vision-language models such as Yang et al. (2023a); Nguyen et al. (2024) due to their limitations in processing long video content as the case of educational videos, although these models demonstrate strong performance on short video clips lasting several seconds. In future work, we plan to enhance the long video understanding capabilities of large multimodal models Zou et al. (2024); Zhou et al. (2024) to enable their applications to VTS.
Evaluation Metrics. We adopt four commonly used metrics, including positive F1 Zhang et al. (2021) (denoted as F1 for brevity), BS@k Gupta et al. (2023), mIoU Mun et al. (2022), and F1@k. Definitions of the four metrics are in Appendix C. Following Gupta et al. (2023), we set k to 30 seconds. We compute the average of these four metrics, denoted by Avg, to measure the overall performance of a model.
@@ 4.2 Results and Analysis
Table 2 compares the performance of baselines (the first group) and variants of our MMVTS models (the second and the third group).
@@
Unimodal performance. For BS@30 on AVLecture, the text-only Longformer (Row 3) outperforms the visual-only BaSSL (Row 2) by a large gain (+25.31), and also surpasses the unsupervised UnsupAVLS by a notable gain (+13.25). Such results are expected since the text modality inherently conveys more precise information for VTS than the vision modality. Notably, the high mIoU of the unsupervised method is attributable to the leakage of the ground-truth number of topics.
@@
Mutimodal performance. As shown in Table 2, Avg (the average of F1, BS@k, mIoU, F1@k) of the multimodal model $SWST_{seq}$ is only comparable to Avg of the text-only LongFormer, suggesting that more data may be necessary to mitigate the discrepancy between pre-training of the language model and fine-tuning with early fusion (as in $SWST_{seq}$), in order to fully exploit the potential of the early fusion strategy. Our MMVTS Baseline1,2,3 simply concatenate unimodal features to predict topic boundaries. Without pre-training on unlabeled data (PT, Eq. 13) and the two auxiliary fine-tuning tasks to enhance multimodal coherence modeling (FT-Coh, Eq. 15), on F1, MMVTS Baseline1 outperforms the text-only Longformer by 2.28 and 2.90 on AVLecture and CLVTS respectively, while on Avg score, MMVTS Baseline1 outperforms Longformer by 1.1 on AVLecture yet slightly underperforms Longformer by 0.65 on CLVTS. These results suggest that simply concatenating unimodal features to predict topic boundaries does not bring consistent gains over unimodal models. The third group in Table 2 evaluates our MMVTS models with the four Multimodal Fusion Layer (MFL) architectures and with pre-training (PT) and fine-tuning (FT-Coh). Overall, after PT and FT-Coh, on AVLecture, our MMVTS model using Co-Attention with MoE as MFLs outperforms all the competitive unsupervised and supervised visual-only and text-only baselines as well as the multimodal $SWST_{seq}$ and achieves the best Avg (4.52 absolute and 7.05% relative gain over previous SOTA), and the bestBS@30 and F1@30 results, setting new SOTA on AVLecture. On CLVTS, our MMVTS model achieves the best F1 (5.21 absolute and 14.98% relative gain over previous SOTA) and yields the Avg score comparable to the best Avg score. On AVLecture, both gains on Avg from our MMVTS model with Co-Attention MoE after PT and FT-Coh over the previous SOTA LongFormercssl and MMVTS Baseline1 are statistically significant ($p<0.05$). Moreover, our MMVTS model significantly outperforms the multimodal model $SWST_{seq}$ on both datasets, demonstrating the effectiveness of our middle fusion strategy and new pre-training and fine-tuning methods. Table 5 shows that our MMVTS model with Co-Attention with MoE has 192M trainable parameters while LongFormercssl has 130M parameters. It is also notable that the performance of models on CLVTS is generally much lower than that on AVLecture, with the best Avg on AVLecture and CLVTS differing by 17.50 (68.84 versus 51.34), indicating a greater challenge to VTS from our CLVTS dataset than the AVLecture dataset.
@@
Comparison with fine-tuning text LLMs. Table 2 also shows that on CLVTS, fine-tuning the powerful pre-trained text LLM Llama-3-8B with our Discrete prompt (Appendix E) achieves the best Avg score 51.34, probably attributable to Llama-3-8B’s extensive pre-training and vast knowledge base; still, the performance of our MMVTS model w/ Merge-Attn and MoE and Co-Attn and MoE after PT and FT-Coh is nearly the same, with Avg 51.33 and 51.27 respectively. However, on AVLecture, fine-tuning Llama-3-8B performs much worse than MMVTS model (60.26 versus our 68.61). Particularly, F1 on both AVLecture and CLVTS from Llama-3-8B are much worse than MMVTS model as we find that Llama-3-8B’s predicted boundaries often have a clip offset. These results underscore the value of our proposed small VTS models, since the efficiency and flexibility of our competitive to superior small models make them indispensable in many real-world applications. Future research could continue exploring how to integrate strengths of LLMs and multimodal approaches for VTS.
Incorporate audio modality. We also explore adding the audio modality (A) to the vision and text modalities (V+T) for our MMVTS model and find that V+T+A slightly improves the Avg score over V+T by 1.09% and 3.32% relatively, as shown in Appendix I. Our ongoing research explores different audio features as well as directly employing visual and audio cues (i.e., V+A) for VTS.
@@
Table 3: Ablation studies of the pre-training tasks on AVLecture test set. The two auxiliary coherence modeling tasks are added in fine-tuning (Eq. 15). For Avg, we report mean and standard deviation from three runs with different random seeds.
We conduct extensive ablation studies to validate effectiveness of the proposed Multimodal Fusion Layers, pre-training and fine-tuning tasks.
@@
(2) Effect of Pre-training tasks. Table 2 shows that for simple concatenation of unimodal features, pre-training before fine-tuning (as Baseline3) outperforms Baseline2 (w/o PT). We conduct ablation studies of the proposed pre-training and fine-tuning tasks on AVLecture. We apply the same fine-tuning with multimodal coherence modeling (FT-Coh, Eq. 15) and compare (a) random initialization of parameters for Multimodal Fusion Layers (w/o pre-training) (b) pre-training the model on unlabeled data. Table 3 shows that w/o pre-training, Co-Attn slightly improves Avg over Merge-Attn by 0.4, and MoE further improves Avg by 0.87. Pre-training improves the performance on all four MFL architectures, with the average Avg score increased by 1.36 (66.29 $\rightarrow$ 67.65). Pre-training also improves model stability as the standard deviations of all w/ PT experiments are less than 1. Table 6 in Appendix further compares the fine-tuning performance after applying different pre-training tasks. Compared to using both pre-training tasks, removing $l_{cma}$ or $l_{vts}$ degrades Avg by 0.4 and 1.64 respectively, indicating that the pretraining task aligned more closely with the downstream task, i.e., $l_{vts}$, yields greater gains.
@@
Table 4: Ablation studies of the two auxiliary coherence modeling fine-tuning tasks on AVLecture test set. Models are initialized from pre-training (Eq. 13).
(3) Effect of Fine-tuning tasks. Table 4 compares different fine-tuning tasks after pre-training. Compared to the standard $l_{vts}$, adding the two auxiliary losses $l_{cma}$ and $l_{mcssl}$ notably improves Avg by 1.12, while decreasing $F_{1}$ by 0.73. These results suggest that while multimodal coherence modeling may slightly compromise the precision of exact matches, it enhances the overall contextual comprehension of a model for VTS. Adding $l_{cma}$ or $l_{mcssl}$ individually improves Avg by 0.53 and 0.57, with improvements mainly on BS@30 and mIoU, suggesting that feature alignment at different granularities may improve fuzzy matching.
@@ 5 Conclusion
We propose a novel supervised VTS model by thoroughly exploring multimodal fusion and coherence modeling. We also introduce a large-scale labeled Chinese Lecture dataset for VTS. Extensive experiments demonstrate the superiority of our model and effectiveness of critical algorithmic designs.
@@
Our MMVTS model leverages the vision modal information encoded by the visual encoder. Considering the computational complexity, we keep the visual encoder frozen while keeping all other parameters trainable. This particular design may result in suboptimal utilization of the extensive multimodal information inherent in the dataset. We plan to further investigate different audio features in multimodal fusion (that is, V+T+A) as well as directly employing visual and audio cues (that is V+A since the current text modality is just ASR 1-best of the audio), to improve the VTS performance. We will continue enhancing the integration of general pre-trained multimodal models and large language models, which may offer a more holistic and effective exploration of multimodal information for VTS. Additionally, we will conduct more explorations and investigate approaches to make Co-Attn with MoE more stable in future work.
Appendix A The CLVTS Dataset
(a) Distribution of Consistency Scores.
(b) Video Duration and Topic Duration (in minutes).
(c) Distribution of Subjects of Lectures.
Figure 2: Statistics of our CLVTS dataset.
A.1 Data Collection and Annotation
The CLVTS dataset is primarily sourced from educational videos, in the form of lectures, from Public Video Platforms [^4] https://study.163.com/ [^5] https://www.bilibili.com/v/knowledge. Specifically, videos are first transcribed by a competitive Automatic Speech Recognition (ASR) system [^6] https://tingwu.aliyun.com/home. We then ask annotators to combine visual and textual (ASR 1-best) information to mark the timestamp (in seconds) at the end of each topic. We ensure high accuracy and reliability of annotations from three aspects, including annotator training, hierarchical topic labeling, and a multi-annotator strategy.
Firstly, before the actual annotation process starts, the annotators take two rounds of training. Each annotator needs to annotate 5 videos in each round; at the end of each round, we assess the annotation quality, provide feedback, and ensure that the annotators address the issues and understand the annotation guideline clearly, at the end of training.
Secondly, during annotation, to help the annotators thoroughly understand the lecture content, we ask the annotators to annotate topics hierarchically, that is, they label both coarse-grained topic (large topic) and fine-grained topic (small topic) boundaries while the large topic boundaries are a subset of the small topic boundaries. We take the small topic boundaries as the final topic boundary labels for supervised VTS modeling.